OpenStudy (anonymous):

Lim n --> Infinity of (-1)^(n/7) I think this has to do with geometric series, but I don't know how

OpenStudy (retireed):

I don't know either, but I am going to graph plot the equation and see what happens as n gets very very large. Can you check my work for me?

OpenStudy (anonymous):

but if you plot graph what you get (-1)^infinity

OpenStudy (anonymous):

what dose it mean in graph

OpenStudy (anonymous):

i think this was has no solution

OpenStudy (anonymous):

what do you say

OpenStudy (anonymous):

and even it happens this question was not exist,

OpenStudy (retireed):

Sorry my OS locked up. I agree with you but that is happen I am trying to figure out.

OpenStudy (anonymous):

but how

OpenStudy (anonymous):

liten bro the quetion asker was gonne how we can know the question was real or from middle of the any solution

OpenStudy (retireed):

Sorry OS issue I had to change computers.

OpenStudy (retireed):

Did you tried it yet.

OpenStudy (retireed):

Yo bro where did you go?

OpenStudy (retireed):

I need to Google limits of negative numbers raised to powers. Bye

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \lim_{n\to\infty}(-1)^{n/7} }\) I am not going to solve this completely, but I will try to give you a good demonstration of this limit. This should be helpful ... Ok, let's go:) \(\large\color{black}{ \large{ \begin{array}{| l | c | r |} \hline ~~~n~~~~ & \displaystyle \lim_{n\to\infty}(-1)^{n/7} \\ \hline ~~~70,000~ & \displaystyle (-1)^{70,000/7} =(-1)^{10,000}=\color{blue}{1} \\ \hline \scr~~~70,007~ & \displaystyle (-1)^{70,007/7} =(-1)^{10,001}=\color{blue}{-1} \\ \hline \scr~~~70,014~ & \scr \displaystyle (-1)^{70,014/7} =(-1)^{10,002}=\color{blue}{1} \\ \hline \scr~~~70,021~ & \scr \displaystyle (-1)^{70,021/7} =(-1)^{10,003}=\color{blue}{-1} \\ \hline \rm so~on~... & \rm~See~the~pattern? \\ \end{array} } }\)

OpenStudy (solomonzelman):

The idea is that you can go infinitely on and on, and your "output" (or the "general term") is not approaching a single value. Remember that for a limit to exist it has to approach ONE value (not multiple values). e.g. A two-sided limit must approach same (and numerical, not \(\infty)\) value from both sides. OR, any one-sided limit must approach one numerical value. If it alternates between +3 or -3 for example, or if it just goes to positive or negative infinity, then it does not exist.

OpenStudy (solomonzelman):

You can easily make sense, just by using the "English" definition. Limit - means that some quantity is given a limit. This means that it is not infinitely large (\(+ \infty\)), or infinitely small (\(- \infty\)). Then it is not "limited", rather, it is unbound, or "u nlimited, if you will. And that also means that there is a single (one) limit to some quantity. (i.e. it doesn't "alternate" - or, doesn't approach many numbers at once.